OpenStudy (anonymous):

What is 0.006 written as a fraction of a percent?

OpenStudy (anonymous):

Move the decimal point two places to the right and then apend the % symbol gives 0.6 %

OpenStudy (anonymous):

0.006= 0.6%= 6/10%= 3/5%